OpenStudy (minime):

5x-3y= 47 6x-y= 7 please solve step by step using substitution

OpenStudy (minime):

ITS A SCHOOL PROJECT HELP PLEASE

OpenStudy (anonymous):

x=-2,y=19

OpenStudy (lalaly):

y=6x-7 substitute that in the first equation 5x-3(6x-7)=47 5x-18x+21=47 -13x=26 x=-2

OpenStudy (lalaly):

now substitute x=-2 in y=6x-7 y=6(-2)-7 = -12-7 = -19